# Kernel Ridge Regression with Predicted Feature Inputs and Applications to Factor-Based Nonparametric Regression

### v1

@@ 3.2 Non-asymptotic upper bounds of the excess risk
In this section we establish non-asymptotic upper bounds of the excess risk $\mathcal{E}(\widehat{f}\circ\widehat{g})$ in (16). An important component of our analysis is the empirical process theory based on local Rademacher complexities developed in Bartlett et al. (2005), building on a line of previous works such as Koltchinskii and Panchenko (2000); Bousquet et al. (2002); Massart (2000); Lugosi and Wegkamp (2004). In our context, local Rademacher complexities translate into the complexity of the RKHS. We begin with a brief review of the basics of RKHS, followed by a discussion of its complexity measures and the assumptions used in our analysis.
In this paper, we make the following assumption on the kernel function. It is satisfied by many widely used kernel functions including the Gaussian and Laplacian kernels.
@@
A key element for analyzing $\widehat{f}$ in (12) is the integral operator $L_{K}:\mathcal{L}^{2}(\rho)\to\mathcal{L}^{2}(\rho)$, defined as
@@
Our next condition assumes that the integral operator $L_{K}$ can be eigen-decomposed.
@@
3 is common in the literature of kernel methods. It is guaranteed, for instance, by the celebrated Mercer’s theorem (Mercer, 1909) when the domain of the kernel function $K$ is compact together with 2, although it can also hold in more general settings (see, e.g., Wainwright (2019, Theorem 12.20)).
Under Assumption 3, the kernel function $K$ admits
@@
$\mathcal{H}_{K}=\Bigg{\{}f=\sum_{j=1}^{\infty}\gamma_{j}\phi_{j}:\ \sum_{j=1}^{\infty}\frac{\gamma_{j}^{2}}{\mu_{j}}<\infty\Bigg{\}}.$
@@
$\langle f_{1},f_{2}\rangle_{K}=\sum_{j=1}^{\infty}\frac{\alpha_{j}\beta_{j}}{\mu_{j}}.$
@@
In classical literature of learning theory, the risk bounds are determined by the richness of the function class. Classical measures include the Vapnik–Chervonenkis dimension (Vapnik and Chervonenkis, 2015), metric entropy (Pollard, 2012) and (local) Rademacher complexity (Koltchinskii and Panchenko, 2000; Koltchinskii, 2001; Mendelson, 2002; Bartlett et al., 2005), see also the references therein. In our case the function class is $\mathcal{H}_{K}$, and its complexity is commonly measured by the following kernel complexity function based on the eigenvalues of the integral operator $L_{K}$,
$R(\delta)=\biggl{(}\frac{1}{n}\sum_{j=1}^{\infty}\min\left\{\delta,{\mu}_{j}\right\}\biggr{)}^{1/2},\qquad\forall\ \delta\geq 0.$ (22)
As detailed in Section 3.4, the (local) Rademacher complexity turns out to be closely related with $R(\delta)$ (Mendelson, 2002). An important quantity in our excess risk bounds is the critical radius $\delta_{n}$, defined as the fixed point to $R(\delta)$, i.e., the positive solution to
$R(\delta)=\delta.$ (23)
It can be verified that $R(\delta)$ is a sub-root function of $\delta$, so that both existence and uniqueness of its fixed point $\delta_{n}$ are guaranteed (Bartlett et al., 2005, Lemma 3.2). Here we recall that
@@
More basic properties of sub-root functions are reviewed in Section C.1.
Finally, we assume the regression error in model (4) to have sub-Gaussian tails. This simplifies our analysis and can be relaxed to sub-exponential tails or bounded finite moments.
@@
Grant model (4) and Assumptions 14. For any $\eta\in(0,1)$, by choosing $\lambda$ in (12) such that
with probability at least $1-\eta$, one has
@@
The explicit dependence of $C$ and $C^{\prime}$ on $\kappa$, $\|f^{*}\|_{\infty}$, $\|f_{\mathcal{H}}\|_{K}$ and $\gamma_{\epsilon}^{2}$ can be found in the proof. Since the proof of Theorem 2 is rather long, we defer its full statement to Section B.2. In Section 3.4 we offer a proof sketch and highlight the main technical difficulties.
As shown in Theorem 2, the excess risk is essentially determined by the critical radius $\delta_{n}$, the kernel-related latent error $\mathbb{E}\Delta_{\widehat{g}}$ and the approximation error $\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$. We discuss each term in detail shortly. It is worth emphasizing that Theorem 2 does not impose any assumptions on the magnitudes of either $\mathbb{E}\Delta_{\widehat{g}}$ or $\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$. Surprisingly, regardless of the magnitude of $\mathbb{E}\Delta_{\widehat{g}}$, it appears additively in the risk bound.
Remark 2 (Comparison with the existing theory of KRR in the classical setting).
Although Theorem 2 is derived for the case where the feature inputs are predicted, it is insightful to compare it with the existing results of KRR in the classical setting. When $Z_{1},\ldots,Z_{n}$ are directly observed and used in (1), setting $\mathbb{E}\Delta_{\widehat{g}}=0$ with $\widehat{g}(z)=z$ for $z\in\mathcal{Z}$ simplifies the risk bound (25) to
@@
If we choose the kernel $K$ using some universal kernels, such as the Gaussian kernel or Laplacian kernel, it is known that the induced RKHS is dense in the space of bounded continuous functions ${\cal C}(\cal Z)$ under the infinity norm (Micchelli et al., 2006). If we further assume $f^{*}\in{\cal C}(\mathcal{Z})$, then $\|f_{\mathcal{H}}-f^{*}\|_{\rho}=0$ for universal kernels. Indeed, for any $\vartheta>0$, there must exist some $f_{\vartheta}\in{\cal H}_{K}$ such that $\|f_{\vartheta}-f^{*}\|_{\infty}\leq\vartheta$. This implies that
@@
The leftmost inequality uses the definition of $f_{\mathcal{H}}$ in 1.
@@ 3.2.1 Discussion on the kernel-related latent error
There exists some positive constant $C_{K}$ such that
@@
$\sup_{z,z^{\prime}\in\mathcal{Z}}\left\|\frac{\partial^{2}K(z,z^{\prime})}{\partial z\partial z^{\prime}}\right\|_{{\rm op}}\leq~{}C.$ (27)
Condition (27) holds for several commonly used kernels, such as the linear kernel and the Gaussian kernel (Blanchard and Zadorozhnyi, 2019; Sun et al., 2022).
Under 5, the kernel-related latent error can be bounded by
@@
entailing consistent prediction of $Z$ itself by using $\widehat{g}(X)$. We remark that this can be relaxed when the kernel function $K$ is invariant to orthogonal transformations, that is, $K(z,z^{\prime})=K(Qz,Qz^{\prime})$ for any $z,z^{\prime}\in\mathcal{Z}$ and any $Q\in\mathbb{O}_{r\times r}$. When this is the case, we only need consistent prediction of $Z$ up to an orthogonal transformation, in the precise sense that
@@
This claim is proved in Section B.4. We give below two common classes of kernel functions that are invariant to orthogonal transformations. Let $h:\mathbb{R}\to\mathbb{R}$ be some kernel-specific function.
@@ 3.2.2 Discussion on the critical radius
The critical radius $\delta_{n}$ defined in (23) is an important factor in the risk bound (25). Finding its exact expression is difficult in general. To quantify its rate of convergence, by the sub-root property of $R(\delta)$ and 2, we prove in Section B.5 the following slow rate of $\delta_{n}$:
$\delta_{n}~{}\leq~{}{\kappa/\sqrt{n}}.$ (30)
Together with Theorem 2, we have the following slow rate of convergence of the excess risk.
@@
Grant model (4) and Assumptions 15. For any $\eta\in(0,1)$ with $\log(1/\eta)\leq\sqrt{n}$, by choosing $\lambda$ as in (24) with $\delta_{n}$ replaced by $\kappa/\sqrt{n}$, with probability at least $1-\eta$, one has
$\mathcal{E}(\widehat{f}\circ\widehat{g})~{}\leq~{}C^{\prime}\left(\frac{\log(1/\eta)}{\sqrt{n}}+\mathbb{E}\|\widehat{g}(X)-Z\|_{2}^{2}+\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}\right).$ (31)
@@
The first $n^{-1/2}$-type rate is well-known in the literature of kernel methods when $Z_{1},\ldots,Z_{n}$ are observed and no assumptions on the smoothness of $f^{*}\in\mathcal{H}_{K}$ or decaying properties of the eigenvalues $\mu_{j}$’s are made. See, for instance, Corollary 5 in Smale and Zhou (2007) and Caponnetto and De Vito (2007). Our new result additionally accounts for both the model misspecification and the error due to predicting feature inputs.
In many situations, however, fast rates of $\delta_{n}$ are obtainable, as we discuss below. For ease of presentation, assume $f_{\mathcal{H}}=f^{*}$ for now. To explain the role of $\delta_{n}$, we follow Yang et al. (2017) by introducing the concept of statistical dimension, which is defined as the last index for which the kernel eigenvalues $\{\mu_{j}\}_{j\geq 1}$ exceed $\delta_{n}$, that is,
$d(\delta_{n})=\max\left\{j\geq 0:{\mu}_{j}\geq\delta_{n}\right\}$
with $\mu_{0}:=\infty$. In view of the expression of $R(\delta)$ in (22), this implies
@@
If $\sum_{j>d(\delta_{n})}\mu_{j}\lesssim d(\delta_{n})\delta_{n}$ holds, we can conclude that $\delta_{n}\asymp{d(\delta_{n})/n}$, hence the name statistical dimension for $d(\delta_{n})$. Any kernel whose eigenvalues satisfy this condition is said to be regular (Yang et al., 2017). More importantly, for any regular kernel, Yang et al. (2017, Theorem 1) shows that in the fixed design setting ($Z_{i}$’s are treated as deterministic), an empirical version of $\delta_{n}$ serves as a fundamental lower bound on the in-sample prediction risk of any estimator:
@@
where the infimum is taken over all estimators based on $\{Y_{i},Z_{i}\}_{i=1}^{n}$, and $\widehat{\delta}_{n}$ is the positive solution to $\widehat{R}(\delta)=\delta$ with $\widehat{R}(\delta)$ being the empirical kernel complexity function, given in (44). In the random design settings, we prove in Theorem 3 that $\widehat{\delta}_{n}$ in the lower bound is indeed replaced by $\delta_{n}$. In the next section we derive explicit expressions of $\delta_{n}$ for some specific kernels.
3.3 Explicit excess risk bounds for specific kernel classes
@@ 3.3.1 Linear kernels
The following corollary states the simplified excess risk bound of Theorem 2 for the linear kernel. Its proof is deferred to Section B.6. The notation $\lesssim_{\eta}$ denotes the inequality holds up to $\log(1/\eta)$.
@@ 3.3.2 Kernels with polynomially decaying eigenvalues
The following corollary states simplified risk bounds of Theorem 2 for such kernels with polynomially decaying eigenvalues. Its proof can be found in Section B.7.
@@ 3.3.3 Kernels with exponentially decaying eigenvalues
The following corollary states the simplified risk bounds of Theorem 2 for such kernels with exponentially decaying eigenvalues. Its proof can be found in Section B.8.
@@ 3.4.1 Bounding the empirical process with a reduction argument
The detailed definition of ${\mathcal{R}_{n}}$ can be found in Section B.2.1. Let $\delta_{x}$ be the fixed point of $\psi_{x}$ such that $\psi_{x}(\delta_{x})=\delta_{x}$. A key step is to establish the following result that holds uniformly over $f\in\mathcal{F}_{b}$,
@@ 3.4.2 Uniformly bounding the cross-term
Here we use $\widehat{\mu}_{x,1}\geq\cdots\geq\widehat{\mu}_{x,n}$ to denote the eigenvalues of the kernel matrix $\mathbf{K}_{x}$ with entries $n^{-1}K(\widehat{g}(X_{i}),\widehat{g}(X_{j}))$ for $i,j\in[n]$. Since $\widehat{R}_{x}(\delta)$ depends only on the predicted inputs, its introduction is pivotal not only for uniformly bounding the cross-term but also, as detailed in the next section, for determining the rate of $\delta_{x}$, the fixed point to the local Rademacher complexity in (38). We show in Lemma 5 of Section B.2.3 that for any $q>0$, with probability $1-\eta$, the following holds uniformly over the function class $\{f\in\mathcal{F}_{b}:\mathbb{E}_{n}[(f\circ\widehat{g})(X)-(f_{\mathcal{H}}\circ\widehat{g})(X)]^{2}\leq q\}$,
@@
so that combining the three displays above yields the final uniform bound for the cross-term, stated in the following lemma, with the detailed proof provided in Section B.2.3.

### v2

@@ 3.2 Non-asymptotic upper bounds of the excess risk
[…]
We make the following assumption on the kernel function, which is satisfied by many commonly used kernels, including the Gaussian and Laplacian.
@@
A key element in our analysis is the integral operator $L_{K}:\mathcal{L}^{2}(\rho)\to\mathcal{L}^{2}(\rho)$, defined as
@@
We assume that the integral operator $L_{K}$ can be eigen-decomposed.
@@
3 is standard in the kernel‐methods literature. It is guaranteed, for example, by Mercer’s theorem (Mercer, 1909) when the kernel domain is compact and 2 holds, though it can also hold in more general settings (see, e.g., Wainwright (2019, Theorem 12.20)). As a result of Assumption 3, the kernel function $K$ admits
@@
$\mathcal{H}_{K}=\{f=\sum_{j=1}^{\infty}\gamma_{j}\phi_{j}:\ \sum_{j=1}^{\infty}{\gamma_{j}^{2}}/{\mu_{j}}<\infty\}.$
@@
$\langle f_{1},f_{2}\rangle_{K}=\sum_{j=1}^{\infty}{\alpha_{j}\beta_{j}}/{\mu_{j}}.$
@@
As mentioned in the Introduction, complexity of the function class $\mathcal{H}_{K}$ is commonly measured by the kernel complexity function $R(\delta)$, given by (7). As detailed in Section 3.4, $R(\delta)$ is related with the (local) Rademacher complexity (Mendelson, 2002). Its fixed point, the positive solution to $R(\delta)=\delta,$ defines the critical radius $\delta_{n}$, a key quantity in our excess risk bounds. Both existence and uniqueness of $\delta_{n}$ are guaranteed as $R(\delta)$ is a sub-root function of $\delta$ (Bartlett et al., 2005, Lemma 3.2). Here we recall that
@@
We review other properties of sub-root functions in Section B.1.
Finally, we assume the regression error in model (4) to be sub-Gaussian. This simplifies our analysis and can be relaxed to sub-exponential tails or bounded finite moments.
@@
Grant model (4) and Assumptions 14. For any $\eta\in(0,1)$, by choosing
$\lambda=C\left(\delta_{n}\log(1/\eta)+\mathbb{E}\Delta_{\widehat{g}}+\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}+\frac{\log(1/\eta)}{n}\right)$ (21)
in (12), with probability at least $1-\eta$, one has
@@
The explicit dependence of $C$ and $C^{\prime}$ on $\kappa$, $\|f^{*}\|_{\infty}$, $\|f_{\mathcal{H}}\|_{K}$ and $\gamma_{\epsilon}^{2}$ can be found in the proof. Since the proof of Theorem 1 is rather long, we offer its sketch and highlight the main technical difficulties in Section 3.4 while defer its full statement to Section A.2.
From Theorem 1, the excess risk is determined by the critical radius $\delta_{n}$, the kernel-related latent error $\mathbb{E}\Delta_{\widehat{g}}$ and the approximation error $\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$. We discuss each term in detail shortly. It is worth emphasizing that Theorem 1 does not impose any assumptions on the magnitudes of either $\mathbb{E}\Delta_{\widehat{g}}$ or $\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$. Surprisingly, regardless of the magnitude of $\mathbb{E}\Delta_{\widehat{g}}$, it appears additively in the risk bound. Finally, achieving the risk bound in (22) requires to choose $\lambda$ as in (12), which in practice could be selected via cross-validation.
Remark 2 (Comparison with the existing theory in the classical setting).
Although Theorem 1 is derived for the case where the feature inputs are predicted, it is informative to compare it with the existing results in the classical setting. When $Z_{1},\ldots,Z_{n}$ are observed and used in (1), setting $\mathbb{E}\Delta_{\widehat{g}}=0$ with $\widehat{g}(z)=z$ for $z\in\mathcal{Z}$ simplifies the risk bound (22) to
@@
If we choose a universal kernel, such as the Gaussian or Laplacian, it is known that the induced RKHS is dense in the space of bounded continuous functions ${\cal C}(\cal Z)$ under the infinity norm (Micchelli et al., 2006). In this case $\|f_{\mathcal{H}}-f^{*}\|_{\rho}=0$ if $f^{*}\in{\cal C}(\mathcal{Z})$. Indeed, for any $\vartheta>0$, there must exist some $f_{\vartheta}\in{\cal H}_{K}$ such that $\|f_{\vartheta}-f^{*}\|_{\infty}\leq\vartheta$, which implies
@@
The leftmost inequality uses the portrayal of $f_{\mathcal{H}}$ in 1.
@@ 3.2.1 Discussion on the kernel-related latent error
There exists some constant $C_{K}>0$ such that
@@
$\sup_{z,z^{\prime}\in\mathcal{Z}}\left\|\frac{\partial^{2}K(z,z^{\prime})}{\partial z\partial z^{\prime}}\right\|_{{\rm op}}\leq~C.$
Such a condition holds for several commonly used kernels, such as the linear kernel and the Gaussian kernel (Blanchard and Zadorozhnyi, 2019; Sun et al., 2022). Under 5, the kernel-related latent error can be bounded by
@@
entailing consistent prediction of $Z$ itself by using $\widehat{g}(X)$. We remark that this can be relaxed when $K$ is invariant to orthogonal transformations, that is, $K(z,z^{\prime})=K(Qz,Qz^{\prime})$ for any $z,z^{\prime}\in\mathcal{Z}$ and any $Q\in\mathbb{O}_{r\times r}$. When this is the case, we only need consistent prediction of $Z$ up to an orthogonal transformation, in the precise sense that (24) can be replaced by
@@
This claim is proved in Section A.4. Two common classes of kernels invariant under orthogonal transformations are
@@
Here $h:\mathbb{R}\to\mathbb{R}$ is a kernel-specific function.
@@ 3.2.2 Discussion on the critical radius
The critical radius $\delta_{n}$, defined as $R(\delta_{n})=\delta_{n}$, is an important factor in the risk bound (22). Finding its exact expression is generally difficult. To quantify its rate, by the sub-root property of $R(\delta)$ and 2, we prove in Section A.5 the slow rate $\delta_{n}\leq{\kappa/\sqrt{n}}.$ Together with Theorem 1, we obtain the following slow rate of the excess risk.
@@
Grant model (4) and Assumptions 15. For any $\eta\in(0,1)$ with $\log(1/\eta)\leq\sqrt{n}$, by choosing $\lambda$ as in (21) with $\delta_{n}$ replaced by $\kappa/\sqrt{n}$,
$\mathbb{P}\left\{\mathcal{E}(\widehat{f}\circ\widehat{g})~\leq~C^{\prime}\left(\frac{\log(1/\eta)}{\sqrt{n}}+\mathbb{E}\|\widehat{g}(X)-Z\|_{2}^{2}+\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}\right)\right\}\geq 1-\eta$ (26)
@@
The first $n^{-1/2}$-type rate is well-known in the literature of kernel methods when $Z_{1},\ldots,Z_{n}$ are observed and no assumptions on the smoothness of $f^{*}\in\mathcal{H}_{K}$ or decaying properties of the eigenvalues $\mu_{j}$’s are made. See, for instance, Corollary 5 in Smale and Zhou (2007). Our new result also accounts for model misspecification and the error of predicting feature inputs.
In many situations, however, fast rates of $\delta_{n}$ are attainable. Yang et al. (2017) introduces the statistical dimension, defined as the largest index $j$ for which the kernel eigenvalues $\mu_{j}$ exceeds $\delta_{n}$, that is, $d(\delta_{n})=\max\left\{j\geq 0:{\mu}_{j}\geq\delta_{n}\right\}$ with $\mu_{0}:=\infty$. By the expression of $R(\delta)$ in (7), one has
@@
If $\sum_{j>d(\delta_{n})}\mu_{j}\lesssim d(\delta_{n})\delta_{n}$ holds, then $\delta_{n}\asymp{d(\delta_{n})/n}$, hence the name statistical dimension for $d(\delta_{n})$. Any kernel whose eigenvalues satisfy such condition is said to be regular (Yang et al., 2017). More importantly, for any regular kernel, Yang et al. (2017, Theorem 1) shows that, in the fixed design setting ($Z_{i}$’s are treated as deterministic) with $f_{\mathcal{H}}=f^{*}$, an empirical version of $\delta_{n}$ serves as a fundamental lower bound for the in-sample prediction risk:
@@
where the infimum is taken over all estimators based on $\{(Y_{i},Z_{i})\}_{i=1}^{n}$, and $\widehat{\delta}_{n}$ is the positive solution to $\widehat{R}(\delta)=\delta$ with $\widehat{R}(\delta)$ being the empirical kernel complexity function, given in (39). For random designs, we show in Theorem 2 of Section 4.3 that $\widehat{\delta}_{n}$ in the lower bound is replaced by $\delta_{n}$. In the next section we derive explicit expressions of $\delta_{n}$ for some specific kernels, with the resulting results summarized in Table 1.
Table 1: Expressions of $\delta_{n}$ for several specific kernels
Linear kernel Polynomial kernels Exponential kernels
$\delta_{n}$ $r/n$ $n^{-\frac{2\alpha}{2\alpha+1}}$ $\log(n)/n$
3.3 Explicit excess risk bounds for specific kernels
@@ 3.3.1 Linear kernels
The following corollary states the simplified excess risk bound of Theorem 1 for the linear kernel. Its proof is deferred to Section A.6. The notation $\lesssim_{\eta}$ denotes the inequality holds up to $\log(1/\eta)$.
@@ 3.3.2 Kernels with polynomially decaying eigenvalues
The following corollary states simplified risk bounds of Theorem 1 for such kernels with polynomially decaying eigenvalues. Its proof can be found in Section A.7.
@@ 3.3.3 Kernels with exponentially decaying eigenvalues
The following corollary states the simplified risk bounds of Theorem 1 for such kernels with exponentially decaying eigenvalues. Its proof can be found in Section A.8.
@@ 3.4.1 Bounding the empirical process with a reduction argument
The detailed definition of $\mathcal{R}_{n}$} can be found in Section A.2.1. Let $\delta_{x}$ be the fixed point of $\psi_{x}$ such that $\psi_{x}(\delta_{x})=\delta_{x}$. A key step is to establish the following result that holds uniformly over $f\in\mathcal{F}_{b}$,
@@ 3.4.2 Uniformly bounding the cross-term
Here we use $\widehat{\mu}_{x,1}\geq\cdots\geq\widehat{\mu}_{x,n}$ to denote the eigenvalues of the kernel matrix $\mathbf{K}_{x}$ with entries $n^{-1}K(\widehat{g}(X_{i}),\widehat{g}(X_{j}))$ for $i,j\in[n]$. Since $\widehat{R}_{x}(\delta)$ depends only on the predicted inputs, its introduction is pivotal not only for uniformly bounding the cross-term but also, as detailed in the next section, for determining the rate of $\delta_{x}$, the fixed point to the local Rademacher complexity in (33). We show in Lemma 5 of Section A.2.3 that for any $q>0$, with probability $1-\eta$, the following holds uniformly over the function class $\{f\in\mathcal{F}_{b}:\mathbb{E}_{n}[(f\circ\widehat{g})(X)-(f_{\mathcal{H}}\circ\widehat{g})(X)]^{2}\leq q\}$,
@@
so that combining the three displays above yields the final uniform bound for the cross-term, stated in the following lemma, with the detailed proof provided in Section A.2.3.